Contribution Margin, Break-Even Units, Contribution Margin Income Statement, Margin of Safety

Zebra Company manufactures custom-designed skins (covers) for iPods® and other portable MP3 devices. Variable costs are $7.70 per custom skin, the price is $14, and fixed costs are $127,890.

Required:

1. What is the contribution margin for one custom skin? Round your answer to the nearest cent.
$ ????? per custom skin

2. How many custom skins must Zebra Company sell to break even?
    ????? custom skins

3. If Zebra Company sells 21,000 custom skins, what is the operating income?
$ ?????

4. Calculate the margin of safety in units and in sales revenue if 21,000 custom skins are sold.

Margin of safety in units ????? units
Margin of safety in sales revenue $ ?????
Answer #1

1. Contribution margin per skin = Selling price - Variable costs

= 14-7.70

= 6.30

Break even units = Fixed cost / Contribution margin per unit

= 127,890 / 6.30

= 20,300

Operating income = (6.30*21,000) - 127,890

= 4410

Margin of Safety units = Sales - Breakeven sales

= 21,000 - 20,300

= 700

Margin of Safety in sales revenue

= 700*14

= 9800